OpenStudy (anonymous):

A student scores 35% and fails by 60 marks while another scores 60% marks and get 40 marks more than the minimum marks required to pass the exam. Find the maximum marks for the examination.

OpenStudy (kropot72):

Let x = minimum pass mark Let y = maximum mark 0.35y = x - 60 ...............(1) 0.60y = x + 40 ..............(2) Now you can solve the system of equations (1) and (2) to find y.

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

y=4 and x=58.6 this is the answer i got how 4 is the maximum mark or 400

OpenStudy (kropot72):

I you subtract equation (1) from eqution (2) the result is 0.25y = 100 ...............(3) Can you solve equation (3) to find y the maximum marks?

OpenStudy (kropot72):

If you....*

OpenStudy (kropot72):

\[y=\frac{100}{0.25}=you\ can\ calculate\]

OpenStudy (anonymous):

ur right.were i was wrong

OpenStudy (kropot72):

You're welcome :)